OpenStudy (anonymous):

MEDAL & FAN :DD The function f(x) = 200 × (1.098)x represents a village's population while it is growing at the rate of 9.8% per year. Create a table to show the village's population at 0, 2, 4, 6, 8, and 10 years from now. Use your table to create a graph that represents the village's population growth. When the population doubles from its current size, the village will need to dig a new water well. To the nearest half of a year, about how long before it is time for the village to dig the new well?

OpenStudy (anonymous):

The function f(x) = 200*(1.098)^x represents a village's population while it is growing at the rate of 9.8% per year Create a table to show the village's population at 0, 2, 4, 6, 8, and 10 years from now. x = 0 f(0) = 200*(1.098)^0 f(0) = 1 x = 2 f(2) = 200*(1.098)^2 f(2) = 241.1208 x = 4 f(4) = 200*(1.098)^4 f(4) = 290.6962 x = 6 f(6) = 350.4645 x = 8 f(8) = 425.5214 x = 10 f(10) = 509.3935 when the population doubles f(x) = 2*200 = 400, lets find x 200*(1.098)^x = 400 by solving we find x = 7.5 years

OpenStudy (anonymous):

@nathshely